OpenStudy (anonymous):

Use the method of undetermined coefficients to find the general solution to: y (double prime)+3y (prime)=2t^(4)+e^(t) sin3t. I got r=0,-3. So the general solution to the homogeneous part is y=c1e^(0t)+c2e^(-3t). I need help finding Y-p

OpenStudy (fibonaccichick666):

first, can I show you how to type that?

OpenStudy (fibonaccichick666):

double primes are just quote marks: " single primes are just 1 : '

OpenStudy (fibonaccichick666):

let me see if I get the same homo answer

OpenStudy (fibonaccichick666):

\[r^2+3r=0\] \[r(r+3)=0\] \[r=0,-3\] check so now, let's talk about, what is the MOUC

OpenStudy (fibonaccichick666):

So, what is the guess for \(t^4\)?

OpenStudy (anonymous):

At^4

OpenStudy (fibonaccichick666):

we remember, it is a polynomial, so we have to include t^4 and all t^n less than t^4, check out that link I attached

OpenStudy (anonymous):

Ok. I'm reading it now. Thanks for the hint of how to type y" and y'

OpenStudy (anonymous):

It should be At^4+Bt^3+Ct^2+Dt. Is it correct?

OpenStudy (fibonaccichick666):

yep, ok so now let's deal with the nasty one

OpenStudy (fibonaccichick666):

e^(t) sin3t.

OpenStudy (anonymous):

e^t cos3t+e^t sin3t. I'm not sure about this one.

OpenStudy (fibonaccichick666):

you forgot constants, make sure you use new letters

OpenStudy (fibonaccichick666):

now just put em together and what do ya got?

OpenStudy (anonymous):

Y (t)=At^4+Bt^3+Ct^2+Dt+K1e^t cos3t+K2e^t sin 3t

OpenStudy (fibonaccichick666):

\[y_p(t)***\]

OpenStudy (fibonaccichick666):

but yes

OpenStudy (anonymous):

So now I should take the first and second derivative, then plug them back into the equation. Right?

OpenStudy (fibonaccichick666):

no,

OpenStudy (fibonaccichick666):

the solution is \[Y=y_c+y_p\] yes?

OpenStudy (anonymous):

Don't I need t solve for A, B, C, D?

OpenStudy (anonymous):

So the answer should be: Y=C1+C2e^-3t+At^4+Bt^3+Ct^2+Dt+K1e^tcos3t+K2e^t sin 3t right?

OpenStudy (fibonaccichick666):

ew this is disgusting, sorry I thought it asked for the particular not the total... ok so yes, you will need to take the entire general solution, differentiate it twice and plug it in

OpenStudy (anonymous):

That's what I thought

OpenStudy (fibonaccichick666):

then just like partial fractions you set the bits equal to each other to determine your coefficients

OpenStudy (anonymous):

Ok. Great. Thanks

OpenStudy (fibonaccichick666):

that is nassstyyyy

OpenStudy (fibonaccichick666):

I recommend variation of parameters to check because that is useless computations

OpenStudy (anonymous):

Yep. Gonna do my best. I have another question if you don't mind.

OpenStudy (fibonaccichick666):

sure just open a new one. I'm working on my analysis hw but I'll look

OpenStudy (anonymous):

Ok. Thanks

OpenStudy (fibonaccichick666):

oh and make sure you tag me